MacDonald had a farm with a certain number of orange trees. He had to cut down 5 trees to control the insects. Each of the remaining trees produced 210 oranges, producing a total harvest of 41,790 oranges.
Write an equation to represent the total harvest in terms of the initial number of trees, t, on MacDonald's farm.

Answer by ankor@dixie-net.com(22740)

MacDonald had a farm with a certain number of orange trees.
He had to cut down 5 trees to control the insects.
Each of the remaining trees produced 210 oranges, producing a total harvest of 41,790 oranges.
Write an equation to represent the total harvest in terms of the initial number of trees, t, on MacDonald's farm.
:
let t = initial number of trees
f(t) = total number of oranges
:
f(t) = 210(t-5) is the equation
:
When f(t) = 41790
210(t-5) = 41790
210t - 1050 = 41790
210t = 41790 + 1050
t = 42840/210
t = 204 trees initially